A Church Songbook

Revised: November 14, 1999

Property of The Church  Last Revised: November 14, 1999  Material used by permission.

1 — What A Mighty God We Serve

Public Domain W&M: Unknown Isaiah 9:6

What a mighty God we serve, What a mighty God we serve, Angels bow before Him, Heaven and earth adore Him, What a mighty God we serve!

1. O Zion, O Zion, that bringest good tidings, Get thee up into the High Mountains Jerusalem, Jerusalem, that bringest good tidings Lift up thy voice with all thy strength Lift it up, be not afraid; Lift it up, be not afraid Say unto the cities of Judah, “Behold your God, Behold your God, Behold your God!”

2 — Finally, My Brethren

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Ephesians 6:10

1. Finally, my brethren, be strong in the Lord, Finally, my brethren, be strong in the Lord, And in the power of His might, And in the power of His might. For it is God at work within you, Both to will and to do, All his glorious pleasure. His truth shall surely endure.

2…. strength …

3…. mercy …

4…. grace …

5. Finally my brethren be strong in the Lord, Finally my brethren be strong in the Lord, And in the power of His might, And in the power of His might, And in the power of His might.

3 — The Lord Of Covenant

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

Lift your hands, for He is Holy. Worship Him in spirit and truth. For His face, is like the sun, The Lord of Covenant.

1. Clap your hands and shout in victory, For the Lamb was slain but lives. Dance for joy and give Him thanks. (Men) For the freedom … (Women) For the freedom His life gives. (Men) For the freedom … (Women) For the freedom His life gives.

2. Come you people, lift your voices, Praise Him with your hearts as one. There is One who turned the key. (Men) And the door … (Women) To His throne is open. (Men) Yes, the door … (Women) To His throne is open.

3. He creates Jerusalem for rejoicing, For rejoicing, And her people for gladness. (Men) He … (Women) He is glad in His people. (Men) Yes, He … (Women) He is glad in His people.

4 — We Will Rejoice

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Song Of Solomon 1:4

We will rejoice in You and be glad, We will extol Your love more than wine Draw me after You and let us run together, We will rejoice in You and be glad

Ch: Lift up the light of Thy countenance, Upon us O Lord Lift up the light of Thy countenance, Upon us O Lord 5 — Praise Him In His Sanctuary

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Psalm 150

1. Praise Him in His sanctuary Praise Him in the skies above Praise Him for the acts of power that He does Praise Him for surpassing greatness With the trumpet, harp and lyre With the tambourines and dancing, praise Him

now

Come and praise Him for the Lord is good And His mercy is everlasting Come and praise Him for the Lord is good And His mercy is everlasting

2. Praise Him with the clashing cymbals Let them hear it far and near With the strings and flutes we’ll praise the Lord

our God Who with majesty is reigning He has power over all Give Him glory and be thankful for His love 6 — Rise Up

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

Rise up, rise up, we are the soldiers of the cross We are the ones who are to glorify the King Creation groans for the Sons of God to come Manifesting all the glory of the King So rise and shine, for your light has come, And the glory of the Lord is upon you Lift up your eyes round about and see All the nations are falling at your feet

Ch: So let’s glorify the Lord, Let’s glorify the King Let’s shout and sing as we take the victory So let’s glorify the Lord, Let’s glorify the King Let’s shout and take the victory!

7 — We Are Marching In Messiah’s Band

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

1. We are marching, in Messiah’s band; The keys of vict’ry in His mighty hand; Let us go on to take the promised land. Raise the anthem, sing the victory song; Praise the Lord for the battle’s won; No weapon formed against Him shall stand.

For the Captain of the Host is Jesus; We are following in His footsteps. No foe can stand against us in the fray.

8 — A Song Of Love

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

A Song of Love, I sing to You A Song of adoration, to the King You are the Lord, the Maker of the universe You are the King of all Kings 9 — Hail To The King

\begin{matrix}c \\ ◯ \end{matrix}1989 Christopher Rath W&M: Christopher Rath

1. I praise You, Lord, for You are my King I praise You, Lord, for You are my King

All hail to the King, All hail to King Jesus All hail to the King of Kings, All hail to the King

2. We praise You, Lord, For You are our King We praise You, Lord, For You are our King

3. I worship You, my Lord and my God I worship You, my Lord and my God

Material used by permission.  Last Revised: November 22, 1993  Property of The Church

10 — The Horse And Rider

\begin{matrix}c \\ ◯ \end{matrix}1950 Mills Music (music only) W&M: Anonymous and I. Miron & J. Grossman Exodus 15:1–2

1. I will sing unto the Lord, for He has triumphed

gloriously The horse and rider thrown into the sea I will sing unto the Lord, for He has triumphed

gloriously The horse and rider thrown into the sea

2. The Lord, my God, my strength, my song, Is now become my victory The Lord, my God, my strength, my song, Is now become my victory

3. The Lord is God and I will praise Him, My Father’s God, and I will exalt Him! The Lord is God and I will praise Him, My Father’s God, and I will exalt Him!

11 — Nothing But The Blood

Public Domain W&M: Robert Lowry Ephesians 1:7

1. What can wash away my sin? Nothing but the blood of Jesus. What can make me whole again? Nothing but the blood of Jesus.

Oh, Precious is the flow That makes me white as snow. No other fount I know, Nothing but the blood of Jesus.

2. This is all my righteousness, Nothing but the blood of Jesus. This is all my hope and peace, Nothing but the blood of Jesus.

12 — Far Above All Rule And Authority

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Ephesians 1:20–21; 2:6

Far above all rule and authority, and power and

dominion Far above all rule and authority, and power and

dominion And every name that is named, not only in this age But also in the one to come And every name that is named, not only in this age But also in the one to come Christ is seated at the Father’s right hand, In heavenly places. We are seated with Him, at the Father’s right hand, In heavenly places.

13 — Sing Hallelujah

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

Sing Hallelujah, sing hallelujah, sing hallelujah, there’s joy in the Lord

1. There is a king in Zion, There is a king in Zion, There is a king in Zion, The City of our God!

2. There’s oil and wine in Zion, There’s oil and wine in Zion, There’s oil and wine in Zion, The City of our God!

3. Let’s go up to Zion, Let’s go up to Zion, Let’s go up to Zion, The City of our God!

4. The church is built on Zion, The church is built on Zion The church is built on Zion, The City of our God!

14 — How Lovely Is Your Dwelling Place

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

How lovely is Your dwelling place, O Lord of Hosts. My soul longs and yearns for Your courts, And my heart and flesh sing for joy to the living God.

One day in Your presence is far better to me than

gold Or to live my whole life somewhere else. And I would rather be a door keeper in Your house Than to take my fate upon myself. You are my Sun and my Shield, You’re my Lover from the start, And the highway to Your city runs through my heart.

15 — Zephaniah 3:17

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

The Lord your God is in your midst, The Lord of lords, who saves, He will exult over you with joy, He will renew you in His love, He will rejoice over you, With shouts of joy, with shouts of joy, With shouts of joy, with shouts of joy, With shouts of joy!

16 — Come And Let Us Go

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Micah 4:2

Come and let us go to the mountain of the Lord, And to the house of our God. Come and let us go to the mountain of the Lord, And to the house of our God. And He will teach us of His ways, And we will walk in His paths. And the law shall go forth from Zion, And the Word of the Lord from Jerusalem.

17 — Behold, I Am The Lord

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

Behold, I am the Lord, the God of all flesh. Is there anything, is there anything, too hard for Me? Is there anything, anything, anything, too hard for

Me? Is there anything, anything, anything, too hard for Me?

18 — Stand Up And Bless

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Nehemiah 9:5

Stand up and bless the Lord your God From everlasting to everlasting Stand up and bless the Lord your God From everlasting to everlasting

Ch: And blessed be your glorious name, O Lord Which is exalted above all blessing and praise And blessed be your glorious name, O Lord Which is exalted, which is exalted 19 — Jesus Is The Lord

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

Jesus is the Lord Jesus the Lord reigns We shall take the kingdoms of this world in His name

Every tribe and nation Every situation Must declare that Jesus is the Lord

Ch: For the Lord our God has delivered Him from

death And established Jesus as Lord He has given Him the power over all that He has

made For our God has made His Christ the Lord 20 — I Will Sing Praises

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

I will sing praises, I will sing praises, I will sing praises to my God I will sing praises, I will sing praises, I will sing praises to my God And when He comes again, The rocks will not have to cry out His people will shout the mighty name Of the King of Glory When the mountains sing forth And all the trees clap their hands His people will shout the name Of the beauty of holiness: Jesus Christ is His name!

21 — Let God Be Magnified

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

Let all those that seek Thee rejoice and be glad In Thee, In Thee And let such as love thy salvation, say Continually, continually, “Let God be magnified, let God be magnified, Let God be magnified.” “Let God be magnified, let God be magnified, Let God be magnified.”

22 — There’s A Light Shining Forth

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

There’s a light shining forth, I can see it on the horizon; It’s the army of God, preparing for war; Coming conquering victorious, O’er the army of satan; Nothing shall stand, before the army of God.

23 — How Great Is Our God

Public Domain W&M: Unknown

How great is our God! How great is His name How great is our God! Forever the same He rolled back the waters, of the mighty Red Sea And He said, “I’ll never leave you, Put your trust in Me!”

24 — Lift Jesus Higher

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown

Lift Jesus higher. Lift Jesus higher. Lift Him up for the world to see. He said, “If I be lifted up from the earth, I will draw all men unto me.”

25 — Garment Of Praise

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Isaiah 61:3

I have put on my garment of praise, I have put on my garment of praise And the spirit of heaviness, is gone from me. I have put on my garment of praise!

26 — I Was Glad

\begin{matrix}c \\ ◯ \end{matrix}Unknown W&M: Unknown Psalm 122:1

I was glad, very glad, when they said unto me Let us go, into the house of the Lord, today There is singing, there is dancing, there is victory In the house of the Lord, today!

27 — Oh, The Blood

Public Domain W&M: Unknown Psalm 51:7

Oh the blood of Jesus. Oh the blood of Jesus. Oh the blood of Jesus. It washes white as snow.